# A Concentration Bound for TD(0) with Function Approximation

### v3

@@ 3 Proof of the Main Result
Let us denote the probability on the right side of the inequality as $p_{n}$. Then
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq p_{n}+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$\leq p_{n-1}+P\left(x_{n-2}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
@@
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq\sum_{m={n_{0}+1}}^{n}p_{m}+P(\|x_{n_{0}}-x^{*}\|>\epsilon).$
The following lemma, proved in Appendix B, gives a bound on the probabilities $p_{m}$:
@@
There exist positive constants $c_{3}$ and $D$ such that for $0<\epsilon\leq 1,0<\delta\leq 1$,
$p_{m}\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)},0<\delta\leq C,$
$p_{m}\leq 2de^{-D\delta/\beta_{n_{0}}(m)},\delta>C,$
where $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$. Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
@@ Appendix A Appendix A: A Martingale Inequality
[…]
@@
$2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$.
@@ B.1 Proof of Lemma 1
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$ $=$ $\|\gamma\sum_{i}\pi(i)\varphi(i)\sum_{j}p(j|i)\varphi(j)^{T}(x-z)$ (13)
$\;\;\;\;\;\;\;-\sum_{i}\pi(i)\varphi(i)\varphi(i)^{T}(x-z)+(x-z)\|^{2}$
@@
Inequality (a) follows from the Cauchy-Schwarz inequality and (b) follows from the observation that $\|Px\|_{D}\leq\|x\|_{D}$, which can be proved as follows.
$\|Px\|^{2}_{D}=\sum_{i\in S}\pi(i)\left(\sum_{j\in S}p(j|i)x_{j}\right)^{2}\leq\sum_{i\in S}\pi(i)\sum_{j\in S}p(j|i)x_{j}^{2}=\sum_{j\in S}\pi(j)x_{j}^{2}=\|x\|^{2}_{D}.$
@@
$\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$
$=\|\Phi^{T}\sqrt{D}(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
@@
$\|\sum_{i}\pi(i)(F(x,i)-F(z,i))\|^{2}$$\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}+\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$ (18)
@@
Note that as the columns of $\Phi$ are linearly independent, $x\neq\theta\implies\Phi x\neq\theta$ and hence $\frac{\|\Phi x\|_{D}}{\|x\|}>0$ when $x\neq\theta$. Also, note that $\min_{x\neq\theta}\frac{\|\Phi x\|_{D}}{\|x\|}=\min_{\|x\|=1}\|\Phi x\|_{D}$ and hence by extreme value theorem, we have that $\min_{\|x\|=1}\|\Phi x\|_{D}$ is attained and is greater than 00. Along with assumption (3), this implies that $\alpha<1$. ∎
@@ B.2 Proof of Lemma 2
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$ (19a)
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}p(j|Y_{k})W(j)\right)x_{k}.$ (19b)
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{j}p(j|Y_{k})U(j)\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)\right)$ (20a)
@@
For (20a), define $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{j}p(j|Y_{k})U(j)$ for $k\geq n_{0}$ and 00 otherwise. This is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k+1})-\sum_{j}(p(j|Y_{k})W(j)\right)x_{k}$ (22a)
@@
Similar to the sequence $\widetilde{U}_{k+1}$, for (22a), define $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{j}p(j|Y_{k})W(j)$ for $k\geq n_{0}$ and 00 otherwise. Note that $\widetilde{W}_{k+1}x_{k}$ is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
@@
$\leq a(n_{0})\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right)$
Before simplifying (22c), we first need to repeat an important simplification from our main proof. Note that for any $0<k\leq m$,
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{i}\pi(i)F(x_{k},i)\right)=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(\widetilde{U}_{k+1}+\widetilde{W}_{k+1}x_{k}\right)+\mu_{m}(n_{0}),$
@@
$\mu_{m}(n_{0})\leq 4a(n_{0})U_{max}+a(n_{0})W_{max}\left(K_{1}+(4+K_{2})\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
@@ B.3 Proof of Lemma 3
We first note that, for $n_{0}\leq k\leq m$, $\|\bar{x}_{k,m}\|\leq\|\bar{x}_{k,m}-\bar{z}_{k,m}\|+\|\bar{z}_{k,m}\|\leq\bar{x}^{\prime}_{m}+\|\bar{z}_{k,m}\|$. By definition,
$\bar{x}^{\prime}_{m}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
@@
$\|\bar{x}_{k,m}\|\leq 1+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}.$
Let $v^{(\ell)}$ denote the $\ell$th component of a vector $v$. Then
@@
$\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\leq c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),$
@@
$C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right),\zeta_{k,m}=\chi(m,k+1)a(k),\varepsilon=1,\gamma_{1}=1.$
Next, we choose suitable $\gamma_{2}$ and $\omega(m)$ such that $\max_{n_{0}\leq k\leq m}\zeta_{k,m}\leq\gamma_{2}\omega(m)$. For this, we use our assumption that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n\geq n_{0}$, to obtain:
$\chi(m,k+1)=\prod_{i=k+1}^{m}(1-a(i))\leq\exp\left(-\sum_{i=k+1}^{m}a(i)\right)\leq\exp\left(-\sum_{i=k+1}^{m}\frac{d_{1}}{i}\right)$
[…]
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\left(\frac{k+1}{k}\right)^{d_{1}}$
$\implies\max_{n_{0}\leq k\leq m}a(k)\chi(m,k+1)\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{k+1}{m}\right)^{d_{1}}\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{2k}{m}\right)^{d_{1}}.$
@@
The factor $d$ comes from the application of union bound to bound the maximum over all dimensions. Under the assumption that $\delta\leq 1$, we have that $e^{-D\delta^{2}/\beta_{n_{0}}(m)}\geq e^{-D\delta/\beta_{n_{0}}(m)}$ and hence $P(\Gamma_{m}\geq\delta)\leq 2de^{D\delta^{2}/\beta_{n_{0}}(m)}$. ∎
@@ B.4 Proof of Corollary 1
$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
@@
[…]

### v4

@@ 4 Proof of the Main Result
Let us denote the probability on the right side of the inequality as $p_{n-1}$. Then
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq p_{n-1}+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$\leq p_{n-2}+P\left(x_{n-2}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)\leq\sum_{m={n_{0}+1}}^{n}p_{m-1}+P(\|x_{n_{0}}-x^{*}\|>\epsilon).$
The probabilities $p_{m}$ can be bounded using standard martingale inequalities as the terms of the martingale difference sequence are almost surely bounded. The following lemma, proved in Appendix B, gives a bound on the probabilities $p_{m}$:
@@
There exists positive constant $D$ such that for $0<\epsilon\leq 1,0<\delta\leq 1$,
$p_{m}\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)}.$
Recall that $d$ here denotes the dimension of the iterates $\{x_{n}\}$.
@@ Appendix A Appendix A: A Martingale Inequality
$E\left[e^{\varepsilon|M_{n}|}\Big|\mathcal{F}_{n-1}\right]\leq C\ \ \forall\;n\geq 1,\mbox{a.s.}$
@@
$2e^{-\frac{D\epsilon}{\omega(n)}},\ \ \mbox{otherwise.}$
@@ B.1 Proof of Lemma 1
$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,i))\|^{2}$ $=$ $\|\gamma\sum_{s\in\mathcal{S}}\pi(s)\varphi(s)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)\varphi(s^{\prime})^{T}(x-z)$ (14)
$\;\;\;\;\;\;\;-\sum_{s\in\mathcal{S}}\pi(s)\varphi(s)\varphi(s)^{T}(x-z)+(x-z)\|^{2}$
@@
Inequality (a) follows from the Cauchy-Schwarz inequality and (b) follows from the observation that $\|Py\|_{D}\leq\|y\|_{D}$, which can be proved as follows.
$\|Py\|^{2}_{D}=\sum_{s\in\mathcal{S}}\pi(s)\left(\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)y(s^{\prime})\right)^{2}\leq\sum_{s\in\mathcal{S}}\pi(s)\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|s)y(s^{\prime})^{2}=\sum_{s^{\prime}\in\mathcal{S}}\pi(s^{\prime})y(s^{\prime})^{2}=\|y\|^{2}_{D}.$
@@
$\|(\gamma\Phi^{T}DP\Phi-\Phi^{T}D\Phi)(x-z)\|^{2}$ $=\|\Phi^{T}\sqrt{D}(\gamma\sqrt{D}P\Phi-\sqrt{D}\Phi)(x-z)\|^{2}$
@@
$\|\sum_{s\in\mathcal{S}}\pi(s)(F(x,s)-F(z,s))\|^{2}$ $\leq$ $\|x-z\|^{2}-2(1-\gamma)\|\Phi(x-z)\|_{D}^{2}+\lambda_{M}^{2}(1+\gamma)^{2}\|\Phi(x-z)\|_{D}^{2}$ (19)
@@
Note that as the columns of $\Phi$ are linearly independent, $x\neq\theta\implies\Phi x\neq\theta$ and hence $\frac{\|\Phi x\|_{D}}{\|x\|}>0$ when $x\neq\theta$. Also, note that $\min_{x\neq\theta}\frac{\|\Phi x\|_{D}}{\|x\|}=\min_{\|x\|=1}\|\Phi x\|_{D}$ and hence by extreme value theorem, we have that $\min_{\|x\|=1}\|\Phi x\|_{D}$ is attained and is greater than 0. Along with assumption (3), this implies that $\alpha<1$. ∎
@@ B.2 Proof of Lemma 2
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$ (20a)
$\;\;\;+\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}.$ (20b)
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(U(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})\right)$ (21a)
@@
For (21a), define $\widetilde{U}_{k+1}=U(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})U(s^{\prime})$ for $k\geq n_{0}$ and 0 otherwise. This is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k})-\sum_{s^{\prime}\in\mathcal{S}}(p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}$
$=\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(W(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}(p(s^{\prime}|Y_{k})W(s^{\prime})\right)x_{k}$ (23a)
@@
Similar to the sequence $\widetilde{U}_{k+1}$, for (23a), define $\widetilde{W}_{k+1}=W(Y_{k+1})-\sum_{s^{\prime}\in\mathcal{S}}p(s^{\prime}|Y_{k})W(s^{\prime})$ for $k\geq n_{0}$ and 0 otherwise. Note that $\widetilde{W}_{k+1}x_{k}$ is a martingale difference sequence with respect to $\{\mathcal{F}_{n}\}$.
@@
$\leq a(n_{0})\left(K_{1}+K_{2}\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right),$
for appropriate $K_{1}$ and $K_{2}$. Before simplifying (23c), we first need to repeat an important simplification from our main proof. Note that for any $0<k\leq m$,
@@
$\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(F(x_{k},Y_{k})-\sum_{s\in\mathcal{S}}\pi(s)F(x_{k},s)\right)$
@@
$\|\mu_{m}(n_{0})\|\leq 4a(n_{0})U_{max}+a(n_{0})W_{max}\left(K_{1}+(4+K_{2})\left(x_{m}^{\prime}+\|x_{n_{0}}-x^{*}\|+\|x^{*}\|\right)\right).$
@@ B.3 Proof of Lemma 3
We first note that, for $n_{0}\leq k\leq m$, $\|\bar{x}_{k,m}\|\leq\|\bar{x}_{k,m}-\bar{z}_{k,m}\|+\|\bar{z}_{k,m}\|\leq\bar{x}^{\prime}_{m,m}+\|\bar{z}_{k,m}\|$. The following follow from the definition of $\mathbb{B}_{m}$. If $\xi\in\mathbb{B}_{m}$, $\bar{x}^{\prime}_{m,m}(\xi)=0$ and if $\xi\notin\mathbb{B}_{m}$, $\bar{x}^{\prime}_{m,m}(\xi)=x_{m}^{\prime}(\xi)\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Hence
$\bar{x}^{\prime}_{m,m}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}.$
@@
$\|\bar{x}_{k,m}\|\leq 1+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}.$
[…]
@@
$\left(M_{k+1}\bar{x}_{k,m}+\widetilde{W}_{k+1}\bar{x}_{k,m}+\widetilde{U}_{k+1}\right)^{(\ell)}\leq c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}\right),$
@@
$C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+c_{1})+1}{1-\alpha-a(n_{0})c_{1}}\right),\zeta_{k,m}=\chi(m,k+1)a(k),\varepsilon=1,\gamma_{1}=1.$
Next, we choose suitable $\gamma_{2}$ and $\omega(m)$ such that $\max_{n_{0}\leq k\leq m}\zeta_{k,m}\leq\gamma_{2}\omega(m)$. For this, we use our assumption that $\frac{d_{1}}{n+1}\leq a(n)\leq d_{3}\left(\frac{1}{n+1}\right)^{d_{2}},\forall\ n\geq n_{0}$, to obtain:
$\chi(m,k+1)=\prod_{i=k+1}^{m}(1-a(i))\leq\exp\left(-\sum_{i=k+1}^{m}a(i)\right)\leq\exp\left(-\sum_{i=k+1}^{m}\frac{d_{1}}{i+1}\right)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\exp\left(-\int_{k+1}^{m+1}\frac{d_{1}}{y+1}dy\right)\leq\exp\left(d_{1}(\log(k+2)-\log(m+2))\right)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\left(\frac{k+2}{m+2}\right)^{d_{1}}$
$\implies\max_{n_{0}\leq k\leq m}a(k)\chi(m,k+1)\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{k+2}{m+2}\right)^{d_{1}}\leq\max_{n_{0}\leq k\leq m}d_{3}\left(\frac{1}{k}\right)^{d_{2}}\left(\frac{2k}{m+2}\right)^{d_{1}}.$
@@
The factor $d$ comes from the application of union bound to bound the maximum over all components. Under the assumption that $\delta\leq 1$, we have that $e^{-D\delta^{2}/\beta_{n_{0}}(m)}\geq e^{-D\delta/\beta_{n_{0}}(m)}$ and hence $P(\Gamma_{m}\geq\delta)\leq 2de^{-D\delta^{2}/\beta_{n_{0}}(m)}$. ∎
@@ B.4 Proof of Corollary 1
$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right),$
@@
$\leq\exp\left(-\int_{n_{0}}^{m}\frac{(1-\alpha)d_{1}}{y+1}dy\right)\leq\exp\left((1-\alpha)d_{1}(\log(k+1)-\log(m+1))\right)$